Looking at the points (3,0) (0,2) on a line (graph) the slope is -2/3 and the value of y when x=2 is ______.
Answer by stanbon(48558)

Looking at the points (3,0) (0,2) on a line (graph) the slope is -2/3 and the value of y when x=2 is ______.
EQUATION: y=(-2/3)x+2
If x=2 y=(-2/3)(2)+2 = -4/3+6/3=2/3
Cheers,
Stan H.